Check whether limit of the function of f(x,y)=4x^5y/x^10+3y^2 exists as (x,y) tends to (0,0).


1
Expert's answer
2022-05-05T10:36:56-0400

If "x=0" then "f(0, y)=\dfrac{0(y)}{0+3y^2}=0." Therefore


"f(x, y)\to 0\text{ as} \ (x,y)\to(0,0) \text{ along the }y-\text{axis}"

For all "x\not=0"


"f(x, x^5)=\dfrac{4x^5(x^5)}{x^{10}+3(x^5)^2}=1\not=0"

"f(x, y)\to 1\not=0\text{ as} \ (x,y)\to(0,0) \text{ along }y=x^5"

Since we have obtained different limits along different paths, limit


"\lim\limits_{(x,y)\to(0,0)}f(x, y)=\lim\limits_{(x,y)\to(0,0)}\dfrac{4x^5y}{x^{10}+3y^2}"

does not exist.